Miller Company's total sales are $210,000. The company's direct labor cost is $25,200, which represents 30% of its total conversion cost and 40 %of its total prime cost. its total selling and administrative expense is $31,500 and its only variable selling and administrative expense is a sales commission of 5 %of sales. The company maintains no beginning or ending inventories and its manufacturing overhead costs are entirely fixed costs.

Required:

1. What is the total manufacturing overhead cost?

2. What is the total direct materials cost?

3. What is the total manufacturing cost?

4. What is the total variable selling and administrative cost?

5. What is the total variable cost?

6. What is the total fixed cost?

7. What is the total contribution margin?
Answer #1

Solution :5 Calculation Of Total Variable Costs :-

Total Variable Costs = Direct Material Costs ( WNote 2)+ Direct Labor Costs + Variable Selling & Administrative Costs (WN3)

= $ 37800 + $ 25200 + $ 10500 = $ 73500

6 Calculation Of Total Fixed Costs :-

Total Fixed Costs = Total Manufacturing Overhead Costs ( WNote 1) + Fixed Selling & Administrative Costs ( WNote 3)

= $ 58800 + $ 21000 = $ 79800

7 Calculation Of Total Contribution Margin :-

Total Contribution Margin = Total Sales Value - Total Variable Costs

= $ 210000 - $ 73500 = $ 136500

Working Note 1 :- Calculation Of Total Manufacturing Overhead Costs

Conversion Costs = Direct Labor Costs + Manufacturing Overhead Costs

Conversion Costs = $ 25200 / 30% = $ 84000

Therefore , Manufacturing Overhead Costs = $ 84000 - $ 25200 = $ 58800 ( All are Fixed Costs)

Working Note 2 :- Calculation Of Total Direct Material Costs

Prime Costs = Direct Material Costs + Direct Labor Costs

Prime Costs = $ 25200 / 40% = $ 63000

Therefore , Direct Material Costs = $ 63000 - $ 25200 = $ 37800

Working Note 3 :- Calculation Of Total Fixed & Variable Selling & Administrative Costs

Total Selling & Administrative Costs = $31500

Variable Selling & Administrative Costs = $21000 * 5% = $ 10500

Therefore , Fixed Selling & Administrative Costs= $ 31500 - $ 10500 = $ 21000